# Energy Efficient Transmission Parameters Selection Method Using Reinforcement Learning in Distributed LoRa Networks

### v1

@@ I Introduction
The Internet of Things (IoT) has gained attention as a mechanism for connecting and controlling objects via the Internet. Consequently, the number of IoT devices is rapidly increasing and is expected to reach 44 billion by 2025 [1]. As a result, communication resource shortages are expected to become more severe. This could lead to network delays and increased power consumption due to channel congestion, interference, and retransmissions. Additionally, it is challenging to recharge a large amount of IoT devices frequently. To address these issues, techniques are needed that can efficiently use wireless resources while consuming low power is necessary.
Wireless communication technologies such as Bluetooth, Zigbee, Wi-Fi, 2G, 3G, and 4G, which have been used for conventional IoT, have limitations such as high power consumption, short transmission distance, and high transmission overhead. On the other hand, low power wide area networking (LPWAN) is gaining attention due to its low power consumption, long-distance communication, low transmission overhead, and low bit rate support [2]. Long Range (LoRa) WAN is one of the representative LPWAN techniques. Compared to other LPWAN techniques, such as SIGFOX, it is open source, which allows the construction of autonomous networks at a low cost without relying on communication infrastructure. In addition, LoRaWAN devices have low power consumption due to infrequent communication, allowing several years of battery life. Additionally, LoRa offers high payload capacity (up to 250 Bytes), and communication ranges from 2 km to 15 km, providing better downlink functionality than Sigfox or Ingenu [2]. Furthermore, LoRa operates in the license-free ISM band, so no radio license is required [3].
@@
On the other hand, distributed methods [7, 8, 9, 10, 11] include the cooperative multi-agent deep reinforcement learning prioritized experience replay (DRL-PER) [7], and the MAB algorithm-based [8, 9, 10, 11, 12] methods. Cooperative multi-agent DRL-PER is a communication resource management approach that uses deep reinforcement learning to achieve high transmission success rates and energy efficiency. However, it requires a large amount of computation and memory, making it unsuitable for resource-constrained IoT devices [7]. The MAB algorithm is a lightweight and can be operated with minimal computational power and memory and exhibits reasonable transmission success rates while maintaining fairness among devices. However, energy efficiency is not considered in the related work [8, 9, 10, 11, 12].
As described above, related work on centralized methods involves the network server acting as an agent to set transmission parameters, which may increase the consumption of communication resources, and they have yet to be verified on real devices [6, 5]. On the other hand, decentralized methods either do not consider energy efficiency or require significant computation and memory, making them difficult to apply to resource-constrained IoT devices [7, 8, 9, 10, 11]. To address these challenges, we propose an implementable transmission parameter selection method that considers energy efficiency for resource-constrained IoT. The main contributions of this paper are summarized as follows.
@@
  • The proposed method is implemented on actual LoRa devices, and its performance is evaluated using the constructed actual massive LoRa network. Experimental results show the superiority of the proposed method in terms of energy efficiency and transmission success rate in the real environment.
This paper is structured as follows: Section II introduces the system model and problem formulation. Section III describes the proposed method. Section IV presents the performance evaluation results. Section V concludes this paper.
@@ II-A System Model
As shown in Fig. 1, this paper considers a LoRa network consisting of one gateway (GW) and $N$ LoRa end devices (EDs). The set of EDs is denoted by $\mathcal{N}$. We define the set of available channels as $\mathcal{D}=\{c_{1},c_{2},\dots,c_{i},\dots,c_{M}\}$ and that of transmission powers as $\mathcal{U}=\{u_{1},u_{2},\dots,u_{i},\dots,u_{P}\}$, where $M$ and $P$ are the number of channels and that of transmission power level, respectively. Each ED selects the combination of channel (CH) and transmission power (TP) using the reinforcement learning algorithm from the parameter set $K=\{k_{1},k_{2},\dots,k_{i},\dots,k_{M+P}\}$, which consists of $M+P$ selections. Each LoRa ED sends data to the GW at regular intervals, performs carrier sense to check if the selected channel is available before transmitting, and transmits data to the GW if the channel is available. When the GW receives data from the ED successfully, it returns an ACK information. If the transmission fails, as with ED3 in Fig. 1, no ACK information is sent from the GW. The transmission parameters selection algorithm is implemented on each LoRa ED to learn to select the transmission parameters using ACK information returned from the GW after each data transmission. This approach also considers energy consumption during data transmission and realizes a distributed reinforcement learning-based transmission parameter selection method.
@@ II-B Energy Consumption Model
    $E_{Active}=E_{WU}+E_{\text{proc}}++E_{\text{ToA}}+E_{\text{R}},$ (1)
@@ II-C Problem Formulation
This paper aims to maximize the energy efficiency of LoRa EDs by using reinforcement learning to select the optimal CH and the smallest transmission power. The object function considered in this paper is as follows.
$(P)=\underset{k_{i}\in K}{\max}\sum_{t=1}^{T}{EE}_{k_{i}}(t),$ (7)
where ${EE}_{k_{i}}(t)$ is defined as:
${EE}_{k_{i}}(t)=\frac{X_{k_{i}}(t)}{E_{Active}},$ (8)
where $X_{k_{i}}(t)$ is the transmission success rate for the transmission parameters combination $k_{i}$ at time $t$, which is defined as:
$X_{k_{i}}(t)=\frac{R_{k_{i}}(t)}{N_{k_{i}}(t)},$ (9)
where $R_{k_{i}}(t)$ and $N_{k_{i}}(t)$ are the number of successful transmissions and that of the selections of parameter set $k_{i}$ at time $t$, respectively. $T$ is the maximum number of transmissions.
@@ III-A What is the MAB Problem?
The MAB problem involves an agent (player) selecting and playing among multiple slot machines [13][14]. The player aims to maximize the reward (coins) obtained by playing repeatedly. The player learns the probability of coins appearing for each slot machine through repeated play. The player must play and explore many slot machines to find the most coin-giving slot machine. However, excessive exploration can prevent maximizing coin acquisition. Thus, the MAB problem is a decision-making problem considering the trade-off between exploring slot machines to find the most coin-giving one and exploiting good slot machines to increase coins.
@@ III-B MAB Algorithm
Auer and Bianchi proposed the UCB series of algorithms [15]. UCB1 is a simple algorithm that balances exploitation and exploration by considering the average reward obtained and the total number of selections made until the present. The UCB1-tuned algorithm enhances the UCB1 algorithm, which considers the variance in the empirical rule of each slot machine. The specific operation involves playing each slot machine once initially and then selecting the machine to play according to the following equation at the $m$-th play.
$l^{*}_{n}=\underset{l_{n}\in L}{\operatorname{argmax}}P_{l_{n}}(m),$ (10)
where $m$ is the number of plays, $L$ is the set of the slot machines, $l_{n}$ is slot machine $n$, $P_{l_{n}}(m)$ is the UCB score of slot machine $n$ at the $m$-th play, which is expressed as follow.
$P_{l_{n}}(m)=\frac{C_{l_{n}}(m)}{S_{l_{n}}(m)}+\sqrt{\frac{\ln m}{S_{l_{n}}(m)}\min\left(\frac{1}{4},V_{l_{n}}(m)\right)},$ (11)
where ${C_{l_{n}}(m)}$ is the total reward of slot machine $n$ at the $m$-th play, ${S_{l_{n}}(m)}$ is the total number of selections of slot machine $n$ at the $m$-th play, $V_{l_{n}}(m)$ is the variance estimator that takes into account the number of plays $m$ and the total number of the selections of slot machine $l_{n}$.
$V_{l_{n}}(m)=\sigma^{2}_{l_{n}}+\sqrt{\frac{2\ln m}{S_{l_{n}}(m)}},$ (12)
where $\sigma_{l_{n}}$ represents the variance of the received rewards of the slot machine $l_{n}$.
@@ III-C Transmission Parameter Selection Method Using the UCB1-Tuned Algorithm
In our proposed method, each LoRa ED, based on the implemented UCB1-tuned algorithm, selects the channel and transmission power. EDs wait for ACK information from the GW and determine rewards based on the presence or absence of this information. The reward for receiving ACK information is defined as 1/$E_{\text{ToA}}$. Otherwise, the reward is defined as 0. Thus, by changing the reward to be given according to the magnitude of power consumption, a smaller transmission power can be selected. Each ED can calculate its energy efficiency after transmission using the selected transmission parameters (channel and transmission power). Each ED starts operation at random timings and repeats the selection of the transmission parameters and transmission using the selected parameters at regular intervals. Based on the description above, to make it easy to understand, we summarize the relationship between the MAB problem and the channel and transmission power selection problem in this work shown in Table I while the proposed method is summarized in Algorithm 1.
TABLE I: Comparison between the MAB Problem and the CH, TP Selection Problem
MAB Problem CH, TP Selection Problem
@@
Objective: Maximize Coins Objective: Optimize Energy Efficiency
@@
0:  $t=0$, $P_{k_{i}}(t)=0$, ${R_{k_{i}}(t)}=0$, ${N_{k_{i}}(t)}=0$, $V_{k_{i}}(t)=0$
1:  for all channel and transmission power combinations do
@@
6:     Select the CH and TP with the maximum UCB score in energy efficiency
7:     Packet transmission use the selected CH and TP
@@
12:  end while
In our proposed method, all variables are initialized as 0 first. Then, each LoRa ED transmits once using each channel and transmission power combination (lines 1-4 in Algorithm 1). After that, the channel and transmission power with the highest value calculated based on (11) are selected, and data are transmitted using the selected parameters (lines 6 and 7 in Algorithm 1). After transmission, the ACK information returned from the GW is checked, and the UCB score is updated based on the presence or absence of ACK information and the level of the selected transmission power (lines 8-10 in Algorithm 1). This operation is repeated a certain number of times, i.e., $T$. Here, $t$ is the number of transmissions, and $P_{k_{i}}(t)$ represents the UCB score of the slot machine $k_{i}$ at the $t$-th play time.
@@ IV-A Experimental Environment and Parameter Settings
In this work, LoRa devices implementing the reinforcement learning algorithm were set up as shown in Fig. 2. The performance evaluations are conducted using this constructed high-density LoRa network. The transmitter and receiver sides consist of the LoRa EDs and the GW, respectively. The transmitters generate and transmit data. The receiver processes the received data to observe the transmission success rate and energy efficiency of each transmitter. Three receivers were set up to imitate GW, and each was assigned a different received channel. The number of LoRa EDs varies from 10 to 30. Each device starts operation at random timings and transmits data every 10 seconds. The bandwidth (BW) and SF are set to 125 kHz and 7, respectively. The number of retransmissions is set to 0. Each LoRa ED selects one channel from five channels, i.e., {920.6, 921.0, 921.4, 921.8, 922.2} MHz, and one transmission parameter from the transmission power set {-3, 1, 5, 9, 13} dBm. The GW can only receive the transmitted data using the following three channels: {921.0, 921.4, 921.8} MHz. The transmission times for each LoRa ED is set to 200. Parameters related to the energy consumption mode used in the performance evaluation are $E_{WU}=56.1$ mW, $E_{proc}=85.8$ mW, $E_{R}=66$ mW, $P_{MCU}=29.7$ mW, $N_{P}=8$ bytes, and $N_{Payload}=36\sim 44$ bytes which depends on the index of the device, the number of transmissions and that of the successful transmission of each LoRa device. The details of the experimental parameters are summarized in Table II.
@@
Selectable CH 920.6, 921.0, 921.4, 921.8, 922.2 MHz
Receivable CH 921.0, 921.4, 921.8 MHz
@@
$E_{WU}$ 56.1 mW
$E_{\text{proc}}$ 85.8 mW
$E_{\text{R}}$ 66 mW
@@ IV-B Comparison Methods
To examine the effectiveness of the proposed method, we compare our proposed method with the $\epsilon$-greedy-based, ADR-Lite, and fixed transmission parameters allocation methods. $\epsilon$-greedy is the simplest MAB algorithm, where LoRa EDs select the combination of the channel and transmission power with the largest reward with probability $1-\epsilon$ and randomly select a combination with probability $\epsilon$. The fixed allocation method pre-assigns CHs evenly to transmitters and transmits at the minimum TP. The ADR-Lite algorithm was introduced as a centralized method, but it is implemented as a distributed method in this performance evaluation. In the ADR-Lite algorithm, the LoRa ED sorts the transmission power in an increased order while the channel is listed according to the channel situation. Specifically, the list of the transmission parameters is set as: {{CH1, -3 dBm}, {CH9, -3 dBm}, {CH3, -3 dBm}, {CH5, -3 dBm}, {CH7, -3 dBm}, {CH1, 1 dBm}, CH9, 1 dBm}, {CH3, 1 dBm}, …, {CH1, 13 dBm}, {CH9, 3 dBm}, {CH3, 13 dBm}, {CH5, 13 dBm}, {CH7, 13 dBm}}, where CH1, CH3, CH5, CH7, and CH9 are the channels with 920.6, 921.0, 921.4, 921.8, and 922.2 MHz, respectively. CH1 and CH9 are unavailable for the receiver, which can be regarded as the channels with the worst situation. The combination of the channel and transmission power located further back of the transmission parameter list, the transmission power is higher while the channel situation is better. In the ADR-Lite algorithm, LoRa ED initiates communication starting with the last combination of the transmission parameters in the list first. If the transmission is successful, the next set of the transmission parameters is halved to the middle value of the first set and the previously selected transmission parameter set in the list; if it fails, the next set of the transmission parameters is set to the transmission parameters in the middle of the last set, and the previously selected transmission parameter set in the list.

### v2

@@ I Introduction
The Internet of Things (IoT) has gained attention as a mechanism for connecting and controlling objects via the Internet. Consequently, the number of IoT devices is rapidly increasing and is expected to reach 44 billion by 2025 [1]. As a result, communication resource shortages are expected to become more severe. This could lead to network delays and increased power consumption due to channel congestion, interference, and retransmissions. Additionally, it is challenging to recharge a large amount of IoT devices frequently. To address these issues, techniques that can efficiently use wireless resources while consuming low power is necessary.
Wireless communication technologies such as Bluetooth, Zigbee, Wi-Fi, and 4G, which have been used for conventional IoT, have limitations such as high power consumption, short transmission distance, and high transmission overhead. On the other hand, Low Power Wide Area Networking (LPWAN) is gaining attention due to its low power consumption, long-distance communication, low transmission overhead, and low bit rate support [2]. Long Range (LoRa) WAN is one of the representative LPWAN techniques. Compared to other LPWAN techniques, such as SIGFOX, it is open source, which allows the construction of autonomous networks at a low cost without relying on communication infrastructure. In addition, LoRa devices have low power consumption due to infrequent communication, allowing several years of battery life. Additionally, LoRa offers high payload capacity (up to 250 Bytes), and communication ranges from 2 km to 15 km, providing better downlink functionality than Sigfox or Ingenu [2]. Furthermore, LoRa operates in the unlicensed ISM band, so no radio license is required [3].
@@
On the other hand, distributed methods [7, 8, 9, 10] include the cooperative multi-agent deep reinforcement learning prioritized experience replay (DRL-PER) [7], and the MAB algorithm-based [8, 9, 10] methods. Cooperative multi-agent DRL-PER is a communication resource management approach that uses deep reinforcement learning to achieve high transmission success rates and energy efficiency. However, it requires a large amount of computation and memory, making it unsuitable for resource-constrained IoT devices [7]. The MAB algorithm is a lightweight and can be operated with minor computational power and memory and exhibits reasonable transmission success rates while maintaining fairness among devices [8, 9, 10, 11, 12, 13]. However, energy efficiency is not considered in the related work [8, 9, 10].
As described above, related work on centralized methods involves the network server acting as an agent to set transmission parameters, which may increase the consumption of communication resources, and they have yet to be verified on real devices [6, 5]. On the other hand, decentralized methods either do not consider energy efficiency or require significant computation and memory, making them difficult to apply to resource-constrained IoT devices [7, 8, 9, 10]. To address these challenges, we propose an implementable transmission parameter selection method that considers energy efficiency for resource-constrained IoT. The main contributions of this paper are summarized as follows.
@@
  • The proposed method is implemented on actual LoRa devices, and its performance is evaluated using the constructed actual high-density LoRa network. Experimental results show the superiority of the proposed method in terms of energy efficiency and transmission success rate in the real environment.
This paper is structured as follows. Section II introduces the system model and problem formulation. Section III describes the proposed method. Section IV presents the performance evaluation. Section V concludes this paper.
@@ II-A System Model
As shown in Fig. 1, this paper considers a LoRa network consisting of one gateway (GW) and $N$ LoRa end devices (EDs). The set of EDs is denoted by $\mathcal{N}$. We define the set of available channels as $\mathcal{C}=\{c_{1},c_{2},\dots,c_{m},\dots,c_{M}\}$ and that of transmission powers as $\mathcal{K}=\{k_{1},k_{2},\dots,k_{p},\dots,k_{P}\}$, where $M$ and $P$ are the number of channels and that of transmission power level, respectively, $c_{m}$ and $k_{p}$ represent the $m$-th channel and the $p$-th transmission power (TP) level, respectively. Each LoRa ED selects one channel from the set $\mathcal{C}$ and one TP level from the set $\mathcal{K}$ before transmitting.
Each LoRa ED sends data to the GW at regular intervals, performs carrier sense to check if the selected channel is available before transmitting, and transmits data to the GW if the channel is available. When the GW receives data from the ED successfully, it returns an ACK information. If the transmission fails, as with ED3 in Fig. 1, no ACK information is sent from the GW. The transmission parameters selection algorithm is implemented on each LoRa ED to learn to select the transmission parameters, e.g., channel and TP level, using ACK information returned from the GW and the power consumption after each time of data transmission.
@@ II-B Energy Consumption Model
    $E_{Active}=E_{WU}+E_{\text{proc}}+E_{\text{ToA}}+E_{\text{R}},$ (1)
@@ II-C Problem Formulation
This paper aims to maximize the energy efficiency of LoRa EDs by using reinforcement learning to select the optimal channel and the smallest TP level. The object function considered in this paper is as follows.
$(P)=\underset{k_{p}\in\mathcal{K},{c_{m}\in\mathcal{C}}}{\max}\sum_{t=1}^{T}{EE}_{(k_{p},c_{m})}(t),$ (7)
where $T$ represents the maximum number of transmission attempts, and ${EE}_{(k_{p},c_{m})}(t)$ is defined as:
${EE}_{(k_{p},c_{m})}(t)=\frac{X_{(k_{p},c_{m})}(t)}{E_{Active}},$ (8)
where $X_{(k_{p},c_{m})}(t)$ is the transmission success rate for the transmission parameters combination $(k_{p},c_{m})$ at time $t$ [11], which is defined as:
$X_{(k_{p},c_{m})}(t)=\frac{R_{(k_{p},c_{m})}(t)}{N_{(k_{p},c_{m})}(t)},$ (9)
where $R_{(k_{p},c_{m})}(t)$ and $N_{(k_{p},c_{m})}(t)$ are the number of successful transmissions and that of the selections of parameter set $(k_{p},c_{m})$ at time $t$, respectively.
@@ III-A What is the MAB Problem?
The MAB problem involves an agent (player) selecting and playing among multiple slot machines [14][15]. The player aims to maximize the reward (coins) obtained by playing repeatedly. The player learns the probability of coins appearing for each slot machine through repeated play. The player should play and explore many slot machines to find the most coin-giving slot machine. However, excessive exploration can prevent maximizing coin acquisition. Thus, the MAB problem is a decision-making problem considering the trade-off between exploring slot machines to find the most coin-giving one and exploiting good slot machines to increase coins.
@@ III-B MAB Algorithm
Auer and Bianchi proposed the UCB series of algorithms [16]. UCB1 is a simple reinforcement learning algorithm that balances exploitation and exploration by considering the average reward obtained and the total number of selections made until the present. The UCB1-tuned algorithm enhances the UCB1 algorithm, which considers the variance in the empirical rule of each slot machine. The specific operation involves playing each slot machine once initially and then selecting the machine to play according to the following equation at the $t$-th play.
$u^{*}_{i}=\underset{u_{i}\in\mathcal{U}}{\operatorname{argmax}}P_{u_{i}}(t),$ (10)
where $P_{u_{i}}(t)$ is the UCB score of slot machine $u_{i}$ at the $t$-th play, which is expressed as follow.
$P_{u_{i}}(t)=\frac{R_{u_{i}}(t)}{N_{u_{i}}(t)}+\sqrt{\frac{\ln t}{N_{u_{i}}(t)}\min\left(\frac{1}{4},V_{u_{i}}(t)\right)},$ (11)
where $R_{u_{i}}(t)$ and $N_{u_{i}}(t)$ are the sum of the rewards and the number of times taken of the slot machine $u_{i}$ prior to time $t$, respectively. $V_{u_{i}}(t)$ is the variance estimator that takes into account the number of plays $t$ and the total number of the selections of slot machine $u_{i}$, which can be expressed as:
$V_{u_{i}}(t)=\sigma^{2}_{u_{i}}+\sqrt{\frac{2\ln t}{N_{u_{i}}(t)}},$ (12)
where $\sigma^{2}_{u_{i}}$ represents the variance of the received rewards of the slot machine $u_{i}$.
@@ III-C Transmission Parameter Selection Method Using the UCB1-Tuned Algorithm
A key feature of the proposed method is its combination of autonomy and lightweight design. Each ED operates autonomously, learning the communication environment and selects appropriate transmission parameters without relying on centralized control. Additionally, because the learning process is based solely on ACK information and TP, the amount of stored information is small. This lightweight design ensures the proposed method can be easily implemented on IoT devices with limited computational and memory resources. Furthermore, each ED learns and prioritizes less congested channels and optimal transmission power levels utilizing ACK information, the channel congestion and communication collisions can be effectively mitigated.
Specifically, in our proposed method, each LoRa ED, based on the implemented UCB1-tuned algorithm, selects the channel and TP level. EDs wait for ACK information from the GW and determine rewards based on the presence or absence of this information. The reward for receiving ACK information is defined as 1/$E_{\text{ToA}}$. Otherwise, the reward is defined as 0. Thus, by changing the reward to be given according to the magnitude of power consumption, a smaller transmission power can be selected. Each ED can calculate its energy efficiency after transmission using the selected transmission parameters (channel and TP level). Each ED starts operation at random timings and repeats the selection of the transmission parameters and transmission using the selected parameters at regular intervals. Based on the description above, to make it easy to understand, we summarize the relationship between the MAB problem and the transmission parameters selection problem in this work shown in Table I while the proposed method is summarized in Algorithm 1.
TABLE I: Comparison between MAB Problem and Transmission Parameters Selection Problem
MAB Problem Channel, TP Selection Problem
@@
Objective: Maximize Coins Objective: Maximize Energy Efficiency
@@
0:  $t=0$, ${R_{(k_{p},c_{m})}(t)}=0$, ${N_{(k_{p},c_{m})}(t)}=0$, $V_{(k_{p},c_{m})}(t)=0$
1:  for all channel and TP combinations do
@@
6:     Select the channel from the channel set $\mathcal{C}$ and TP level form the TP set $\mathcal{K}$ with the maximum UCB score in energy efficiency, respectively.
7:     Packet transmission use the selected channel and TP
@@
12:     Sleep mode
13:  end while
In the proposed method, all variables are initialized as 0 first. Then, each LoRa ED transmits once using each channel and TP level combination (lines 1-4 in Algorithm 1). After that, the channel and TP level with the highest value calculated based on (11) are selected, and data are transmitted using the selected parameters (lines 6 and 7 in Algorithm 1). After transmission, the ACK information returned from the GW is checked, and the UCB score is updated based on the presence or absence of ACK information and the level of the selected TP (lines 8-10 in Algorithm 1). This operation is repeated a certain number of times, i.e., $T$. Here, $t$ is the number of transmissions.
@@ IV-A Experimental Environment and Parameter Settings
In this work, LoRa EDs implementing the proposed method were set up as shown in Fig. 2. The transmitter and receiver sides consist of the LoRa EDs and the GW, respectively. The transmitters generate and transmit data. The receiver processes the received data to observe the transmission success rate and energy efficiency of each transmitter. Three receivers were set up to imitate GW, and each was assigned a different received channel. The number of LoRa EDs varies from 10 to 30 in the performance evaluations. Each device starts operation at random timings and transmits data every 10 seconds. The bandwidth (BW) and SF are set to 125 kHz and 7, respectively. The number of retransmissions is set to 0. Each LoRa ED selects one channel from five channels, i.e., {920.6, 921.0, 921.4, 921.8, 922.2} MHz, and one transmission power from the transmission power set {-3, 1, 5, 9, 13} dBm. The GW can only receive the transmitted data using the following three channels: {921.0, 921.4, 921.8} MHz. The transmission times for each LoRa ED is set to 200. Parameters related to the energy consumption mode used in the performance evaluation are $E_{WU}=(56.1*T_{\text{WU}})$ mWh, $E_{proc}=(85.8*T_{\text{proc}})$ mWh, $E_{R}=(66*T_{\text{R}}$) mWh, $P_{MCU}=29.7$ mW, $N_{P}=8$bytes, $N_{Payload}=36\sim 44$ bytes. Here, $T_{\text{WU}}$ represents the wake-up time of the LoRa device, $T_{\text{proc}}$ is the processing time for selecting transmission parameters by the microcontroller, and $T_{\text{R}}$ is the reception time of the device. The details of the experimental parameters are summarized in Table II.
@@
Selectable channel 920.6, 921.0, 921.4, 921.8, 922.2 MHz
Receivable channel 921.0, 921.4, 921.8 MHz
@@
$E_{WU}$ $56.1*T_{\text{WU}}$ mWh
$E_{\text{proc}}$ $85.8*T_{\text{proc}}$ mWh
$E_{\text{R}}$ $66*T_{\text{R}}$ mWh
@@ IV-B Comparison Methods
To examine the effectiveness of the proposed method, we compare our proposed method with the $\epsilon$-greedy-based, ADR-Lite, and fixed transmission parameters allocation methods. $\epsilon$-greedy is the simplest MAB algorithm, where LoRa EDs select the combination of the channel and transmission power with the largest reward with probability $1-\epsilon$ and randomly select a combination with probability $\epsilon$. The fixed allocation method pre-assigns CHs evenly to transmitters and transmits at the minimum TP. The ADR-Lite algorithm was introduced as a centralized method, but it is implemented as a distributed method in this performance evaluation. In the ADR-Lite algorithm, the LoRa ED sorts the transmission power in an increased order while the channel is listed according to the channel situation. Specifically, the list of the transmission parameters is set as: {{CH1, -3 dBm}, {CH9, -3 dBm}, {CH3, -3 dBm}, {CH5, -3 dBm}, {CH7, -3 dBm}, {CH1, 1 dBm}, CH9, 1 dBm}, {CH3, 1 dBm}, …, {CH1, 13 dBm}, {CH9, 3 dBm}, {CH3, 13 dBm}, {CH5, 13 dBm}, {CH7, 13 dBm}}, where CH1, CH3, CH5, CH7, and CH9 are the channels with 920.6, 921.0, 921.4, 921.8, and 922.2 MHz, respectively. CH1 and CH9 are unavailable for the receiver, which can be regarded as the channels with the worst situation. The combination of the channel and TP located further back of the transmission parameter list, the TP is higher while the channel situation is better. In the ADR-Lite algorithm, LoRa ED initiates communication starting with the last combination of the transmission parameters in the list first. If the transmission is successful, the next set of the transmission parameters is halved to the middle value of the first set and the previously selected transmission parameter set in the list; if it fails, the next set of the transmission parameters is set to the transmission parameters in the middle of the last set, and the previously selected transmission parameter set in the list.